Suppose 9.1 g of octane is mixed with 59.5 g of oxygen. Calculate the minimum mass of octane that could be leftover by the chemical reaction
1
Expert's answer
2020-05-15T10:23:11-0400

C8H18 + 12.5O2 = 8CO2 + 9H2O

n(C8H18) = "{\frac {9.1}{114}}=0.08" mol

n(O2) = "{\frac {59.5}{32}}=1.86" mol

the oxygen is taken in excess, so octane itself is a limiting reactant, for 0.08 moles of octane 1 mol of oxygen is needed for complete combustion. So all octane will be burnt, and 0.86 mol of oxygen is left.